Elin Saunders, of Lawn Pro Company, designs and installs custom lawn and garden irrigation systems for homes and businesses throughout the province. Each job is different, requiring different materials and labour for installing the systems. Lawn Pro estimated the following for the year:

    Number of installations            340

    Number of direct labour hours        5 440

    Direct materials cost                R102 000

    Direct labour cost                R54 400

    Overhead cost                    R59 840

During the year, the following actual amounts were experienced:

Number of installations            290

    Number of direct labour hours        4 350

    Direct materials cost                R87 000

    Direct labour cost                R43 500

    Overhead cost                    R53 460

Required:

2.1    Should Lawn Pro use process costing or job costing? Explain.

2.2    If Lawn Pro uses a normal costing system and overhead is applied on the basis of direct labour hours, calculate the cost on an installation that takes R3 500 of direct materials and 30 direct labour hours.

2.3    Explain why Lawn Pro would have difficulty using an actual costing system.
Answer #1

Since Lawn Pro designs custom products and each job is different requiring different material and labour it should use job costing Method. Here there is no transfer of costs from one process to another . And costs are calculated for each job.

Under a normal costing system the cost of installation would be

Material cost R3500

Labour cost

(54400/5440*30) R300

overhead cost at a predetermined overhead rate =30*11 R330

(predetermined overhead rate =estimated overhead/estimated labour hours)

=59840/5440=11 per labour hour

total cost R 4130

Since the number of installations in every job is different. Lawn Pro will face difficulty in allocating overheads to every job hence it cannot use actual costing method instead job costing method estimating overhead cost will have to be used